OpenStudy (anonymous):

A 20 g bullet inelastically collides with a 3 kg wooden block at rest. After the collision, the bullet-block system moves with a velocity of 1.5 m/s. How much energy is lost in the collision?______J I got 563.6227 J. is this correct? Thanks guys!

OpenStudy (theeric):

I got under 500J..... I can recalculate!

OpenStudy (theeric):

I found a mistake. I did it two ways, and now they both get the same result. I got less than your answer, greater than 500J. How did you approach this?

OpenStudy (anonymous):

I calculated the momentum after and then set it equal to what it would be before the collision to find the initial velocity. \[P_b=.002v_i+0 P_b=.002v_i \] \[P_A=.002v_i=1.002(1.5m/s)\] \[v_i=751.5m/s\] \[k_b=.5(.002)(751.5)^2\] \[K_b=564J\] \[K_A=.5(1.002)(1.5)^2\] \[K_A=1.12725\] \[K_b-K_A=56362275J\]

OpenStudy (anonymous):

ahhh I found my mistake... I think!

OpenStudy (theeric):

Is it in your \(P_A\)?

OpenStudy (theeric):

Or, rather the value for the bullet mass in general!

OpenStudy (anonymous):

okay with my second go round.. ( I see that my decimal has been off for the weight=.02kg) my new answer is 57.375 J

OpenStudy (theeric):

So, \(P_b\) is good, but still has an unknown. \(P_A\) looks to consider that the mass of the block is \(1\rm\ kg\). Did you fix this in your new solution?

OpenStudy (anonymous):

I thought that the momentum would be the same before and after so that is why I set \[P_A=P_B\] that is how i found the initial v.

OpenStudy (theeric):

Yup, that's right! But at first you wrote \(P_A=.002v_i=1.002(1.5m/s)\) and then you realized you were off by a decimal place and so \(P_A=.02v_i=1.02(1.5m/s)\) I think that it should be \(P_A=.02[{\rm kg}]v_i=(\color{#1111DD}3.02[{\rm kg}])~~(1.5[{\rm m/s}])\)

OpenStudy (anonymous):

omg..I can't believe I've been using a completely different number! Thanks a lot!

OpenStudy (theeric):

Haha, no problem! I've been there.

OpenStudy (anonymous):

okay... 509.625J final answer lol

OpenStudy (theeric):

I got the same result :)

OpenStudy (anonymous):

Yay!

OpenStudy (theeric):

Congrats! I think we got it correct! Let me know if you learn otherwise...

OpenStudy (anonymous):

surely will!

OpenStudy (theeric):

Alright! Take care!

OpenStudy (anonymous):

@theEric it was right!!

OpenStudy (theeric):

\(\huge\color{#1144FF}{\large\ \ o\ o\ \\\smile}\)\(~\\\large\quad \it\color{green}{awesome!}\)